### v4

# V-RoAst: Visual Road Assessment Can VLM be a Road Safety Assessor Using the iRAP Standard?

[…]
V-RoAst: Visual Road Assessment
Can VLM be a Road Safety Assessor Using the iRAP Standard?
Natchapon Jongwiriyanurak1,4, Zichao Zeng1,4,5, June Moh Goo1,5, Xinglei Wang1,4, Ilya Ilyankou1,4,
Kerkritt Sriroongvikrai2, Nicola Christie1, Meihui Wang1,4, Huanfa Chen3, James Haworth1,4
1 Department of Civil, Environmental and Geomatic Engineering (CEGE), UCL
2 Department of Civil Engineering, Chulalongkorn University
3 Centre for Advanced Spatial Analysis (CASA), UCL 4 SpaceTimeLab 5 3DIMPACT
Corresponding author: natchapon.jongwiriyanurak.20@ucl.ac.ukCorresponding author: j.haworth@ucl.ac.uk
Road safety assessments are critical yet costly, especially in Low- and Middle-Income Countries (LMICs), where most roads remain unrated. Traditional methods require expert annotation and training data, while supervised learning-based approaches struggle to generalise across regions. In this paper, we introduce V-RoAst, a zero-shot Visual Question Answering (VQA) framework using Vision-Language Models (VLMs) to classify road safety attributes defined by the iRAP standard. We introduce the first open-source dataset from ThaiRAP, consisting of over 2,000 curated street-level images from Thailand annotated for this task. We evaluate Gemini-1.5-flash and GPT-4o-mini on this dataset and benchmark their performance against VGGNet and ResNet baselines. While VLMs underperform on spatial awareness, they generalise well to unseen classes and offer flexible prompt-based reasoning without retraining. Our results show that VLMs can serve as automatic road assessment tools when integrated with complementary data. This work is the first to explore VLMs for zero-shot infrastructure risk assessment and opens new directions for automatic, low-cost road safety mapping. Code and dataset: https://github.com/PongNJ/V-RoAst.
1 Introduction
Figure 1: Locations of the ThaiRAP dataset and an example of images in a road segment. The numbers on the images indicate their order within the section.
Road crashes are a major contributor to global mortality, disproportionately affecting Low- and Middle-Income Countries (LMICs) [43, 7]. Scalable, low-cost assessment methods are crucial to meeting road safety targets under the UN’s Global Plan for the Decade of Action aiming to ensure that all new roads are built to achieve a rating of at least 3 stars according to the International Road Assessment Programme (iRAP) standard, which rates roads on a scale from 1 to 5, with 5 indicating a safe road and 1 indicating an unsafe road [43]. Furthermore, another objective is to improve 75% of existing roads to more than 3 stars by 2030.
The workflow of iRAP involves road surveys, coding attributes, developing the model, and analysing the results. From iRAP survey manual [^1] https://resources.irap.org/Specifications/iRAP_Survey_Manual.pdf, the process requires vehicles and sensors to capture accurate georeferenced images for coding. Once georeferenced images are obtained, trained coders must examine and classify the images according to the codebook manual [^2] https://resources.irap.org/Specifications/iRAP_Coding_Manual_Drive _on_Right.pdf, which requires training and experience. Currently, the road ratings cover mostly highways, as the high cost of surveys has limited broader coverage. Importantly, it is almost impossible for LMICs to assess all roads following this standard. This leaves the vast majority of the road network unrated, making it difficult to reveal the infrastructure risk factors contributing to road deaths.
To reduce the cost of manual road assessments, automated detection of road features from imagery has become a widely adopted approach—typically using Convolutional Neural Networks (CNNs). While more affordable than human labelling, these models require task-specific training data, making them time-consuming to scale across regions due to visual variability in road environments across cities and countries [23, 34, 21, 38, 37]. Other studies have explored alternative data sources to overcome these limitations, including LiDAR [5], satellite imagery [6, 1], UAV imagery [4], and GPS traces [46].
Recent studies have explored Visual Language Models (VLMs) for various vision tasks, leveraging their ability to perform without additional training due to large-scale image-text pretraining. Techniques like prompt engineering, Retrieval-Augmented Generation (RAG) [26], and fine-tuning [44] have been proposed to enhance their performance. VLMs have shown promise in zero-shot tasks such as building age classification [48], landscape image tagging [19], and motorcycle risk assessment [22]. However, their potential for iRAP-based road attribute classification remains unexplored.
This work investigates VLMs for supporting road assessments by developing prompts to classify iRAP attributes by mimicking a coder observing an image and categorising the attributes, as described in the codebook manual. Additionally, we evaluate the feasibility of leveraging the zero-shot capabilities of VLMs for this task, assuming the models are sufficiently advanced for practical application. In summary, we state the main contributions of our work as follows:
  • Open-Source VLM Benchmark: We propose a new image classification task for VLMs with a real-world open-source dataset from ThaiRAP.
  • Prompt Optimisation: We optimise the prompts and evaluate the potential of using VLMs to code road attributes using Gemini-1.5-Flash and GPT-4o-mini compared to traditional computer vision models.
  • Zero-Shot Classification: We demonstrate that VLMs can classify iRAP attributes in a zero-shot setting, achieving competitive performance without any task-specific training. VLMs excel at generalising to unseen images and offer strong potential for further improvement through in-context learning or lightweight fine-tuning.
  • Automatic Road Assessment Framework: We present an automatic approach using crowdsourced imagery from Mapillary to estimate star ratings.
2 Literature review
Automated Road Attribute Classification
Computer vision has long been used to support road safety analysis, particularly in detecting surface defects (e.g., cracks, potholes) and classifying road features [21, 18, 12, 30]. Several studies have attempted to automate iRAP attribute extraction using CNN-based models trained on labelled street-level imagery [23, 38, 40]. These approaches, while effective in structured environments, require extensive training data and struggle to generalise across regions due to visual variation and data imbalance [14].
To reduce annotation effort, multi-task learning frameworks have been applied to simultaneously predict several attributes [24, 23]. However, such models often require domain adaptation techniques to maintain performance across geographic regions [3, 28, 25] and are limited to predefined label sets [24]. Most can fail to scale to under-resourced contexts, where training data is scarce or unavailable [24].
Visual Language Models
Recent advancements in large Vision-Language Models (VLMs) such as Gemini [10], GPT-4o [33], and LLaVA [29] have shown promise in zero-shot classification, captioning, and visual reasoning. These models are pretrained on large-scale image-text pairs and can generate responses conditioned on natural language prompts, enabling adaptation to new tasks without fine-tuning.
VLMs have been applied to various spatial tasks, including building understanding and captioning [11, 27], and autonomous driving [42, 20, 49, 45]. While their use in urban analytics is growing, applications in policy-driven, structured classification tasks—such as iRAP-based road assessments—remain largely unexplored.
Visual Question Answering (VQA)
VQA was initially introduced as a new task in the computer vision and natural language processing domains [2]. The task involves answering open-ended questions based on images. Several datasets are commonly used to evaluate models, including GQA [17], OK-VQA [31], A-OKVQA [31], and MMMU [15].
In addition, VQA datasets focusing on autonomous vehicles, such as KITTI [9] and NuScenes-QA [35], have been used to advance the field. Jain et al. [20] found that GPT-4 performs well in robust driving scenarios that require semantic understanding. Similarly, Dihan et al. [8] reported that Gemini‑1.5‑Pro demonstrated competitive performance on map-based geospatial reasoning tasks, highlighting its potential in urban analytics applications.
This work utilises a real-world dataset to construct a new image classification benchmark specifically designed for VLMs. Rather than training a new VLM, we explore the potential of existing general-purpose VLMs (e.g., Gemini and GPT) to serve as road assessment assessors in a zero-shot setting. Although non-open-source models incur some usage costs, our goal is to demonstrate that road assessment tasks can benefit from the reasoning capabilities of existing VLMs. This setup also opens possibilities for prompt engineering, RAG, and lightweight fine-tuning to enhance performance.
3 Dataset Construction
3.1 Data Collection
We provide a real-world iRAP-compliant road assessment dataset comprising 2,037 street-level images (1600×1200 pixels) captured across Bangkok, Pathum Thani, and Phra Nakhon Si Ayutthaya. The images represent 519 road segments, with 1–4 images per 100 m segment, and are annotated by trained iRAP coders. Figure 1 shows the geographic distribution. Unlike most iRAP datasets, this dataset is publicly available to support research in automated road safety assessment.
3.2 Dataset Annotation
Figure 2: V-RoAst Dataset Annotation Process
Figure 3: Code distribution: the numbers at the top indicate the unique codes (representing all possible codes). The following abbreviations are used: (OF) = Observed Flow, (F) = Facilities, (DS) = Driver-Side, (PS) = Passenger-Side, (PC) = Pedestrian Crossing, (POF) = Pedestrian Observed Flow, (RS) = Roadside Severity, (SZ) = School Zone.
Each road segment is labelled with 52 iRAP-defined attributes, including number of lanes, road curvature, roadside hazards, and pedestrian facilities. Labels were assigned by trained road safety engineers, following iRAP’s global coding protocol. Figure 2 illustrates the annotation process, and Figure 3 shows class distribution across all coded attributes. The dataset includes both categorical and ordinal attributes, enabling fine-grained evaluation of classification performance across imbalanced classes.
3.3 Dataset Statistics
The dataset includes ordinal attributes, enabling fine-grained evaluation of classification models. However, the number of classes per attribute varies (shown at the top of each bar in Figure 3). The dataset exhibits strong class imbalance: 11 attributes contain only a single class, making them unusable for supervised training, while several others include rare classes with fewer than 10 samples.
Not all attributes span the full set of possible iRAP codes(classes). For example, “Bicycle Observed Flow” contains only 6 of the official classes in this dataset, and “Centreline Rumble Strips” appears with just 2. These limitations highlight the practical constraints of real-world data collection and motivate the use of zero-shot classification approaches that do not require balanced training data.
3.4 Data Preprocessing
To compare our method with traditional computer vision baselines, the dataset (n = 2,037) was divided into training (1,274 original + 464 augmented), testing (492), validation (243), and unseen (28) sets. We ensured class-balanced splits for each attribute where possible. The splitting followed these rules:
  1. 1.
    Attributes with only one class were excluded.
  2. 2.
    Classes with 5–11 samples were augmented.
  3. 3.
    Classes with 4 or fewer samples were augmented only if the attribute had two classes.
  4. 4.
    If an attribute had more than two classes and a class contained $\leq$ 4 samples, those samples were moved to the unseen set.
To mitigate class imbalance (Rules 2 and 3), five types of noise were applied to selected images: Gaussian, salt-and-pepper, speckle, periodic, and quantisation noise. All augmented images were included only in the training set.
The unseen set was reserved for evaluating zero-shot prediction performance. Baseline models were trained using the training set and validated on the validation set. Final evaluations for both baselines and VLMs were conducted on the testing and unseen sets.
4 Proposed Method
4.1 Problem Definition
Given an input consisting image $I$ and associated metadata $M$ including image ID, latitude, and longitude, the objective is to accurately classify it across 52 distinct attribute types, where each attribute type is treated as a separate multi-class classification problem. The Visual Road Assessment (V-RoAst) framework incorporates a vision-language model $f$, guided by a system prompt $T_{S}$ and a user prompt $T_{U}$, to perform the classification. The output is a set of 52 predictions $\mathcal{A}=\{a_{1},a_{2},...a_{52}\}$, where $\mathcal{A}$ denotes the set of predicted attributes, $a_{i}\in C_{i}$, and $C_{i}$ is the set of possible classes for the $i$-th attribute type.
For each road segment, we are provided with between 1 and 4 images, each associated with its own set of predicted attributes. The goal is to derive a single aggregated attribute set per road that reflects the most critical conditions, consistent with road safety assessment standards. Specifically, for each attribute type, each image produces a prediction $a_{i}^{j}\in\mathcal{A}$, where $j\in\{1,...,n\}(n\leq 4)$ indexes the images for a given road. The predicted attribute $\hat{a}_{i}^{r_{k}}$ for the road $r_{k}$ is selected as the highest-risk class among the image-level predictions, according to a predefined risk ranking from the iRAP specification. The aggregated road-level preditions $\hat{\mathcal{A}}^{r_{k}}=\{\hat{a}_{1}^{r_{k}},...,\hat{a}_{52}^{r_{k}}\}$ and then compared against the ground truth road attributes $\hat{\mathcal{A}}_{g}^{r_{k}}=\{\hat{a}_{g1}^{r_{k}},...,\hat{a}_{g52}^{r_{k}}\}$ using standard multi-class classification metrics.
4.2 Visual Road Assessment Framework
Figure 4: Framework of V-RoAst for Visual Road Assessment
Figure 5: System Prompt from Figure 4
V-RoAst is a framework designed to classify 52 iRAP attributes $\mathcal{A}$ from the input image $I$, as shown in Figure 4. This framework is easily applicable in any city, requiring minimal expertise. The framework utilises VLMs via text input for system instructions $T_{S}$ and user prompts $T_{U}$ which include $I$ and $M$. The workflow operates by inputting an image $I$ and its associated metadata $M$, including task specifications, local context, attribute details, and output format into the VLM $f$. This process enables the generation of attribute-level predictions $\mathcal{A}$ for each image $I$.
4.2.1 System Instruction
As shown in Figure 5, System Instruction ($T_{S}$) is divided into 4 parts, including task specification, local context, attribute details, and output format as suggested by OpenAI [32] and Gemini [10] technical report.
  1. 1.
    Task Specification: This component provides step-by-step instructions for processing each image. The model is guided to classify each <<<Attribute>>> by selecting the most appropriate <<<Category class>>> based on its visual description. This section also specifies how the output should be formatted. Notably, {country} is included to localise the analysis for each region.
  2. 2.
    Local Context: Information such as which side of the road vehicles drive on is provided through {local_context}. This helps the VLMs interpret images more accurately by incorporating local traffic rules and infrastructure norms. It is especially useful for visually ambiguous attributes—like the presence of service roads—where local knowledge improves classification. Including this input mimics how human coders use familiarity with the area to make more accurate assessments.
  3. 3.
    Attribute Details: Each attribute is defined with its name, description ($<$Attribute description$>$), and possible classes ($<$Category class$>$), accompanied by descriptions of each category ($<$Category description$>$). These structured definitions ensure that the models have clear guidelines for classification. Note that this section includes approximately 100,000 characters (or 20,000 tokens) from the iRAP coding manual, which this work made available in JSON format.
  4. 4.
    Output Format: The results are returned JSON format to maintain consistency. The models are instructed to select the best match from the provided $<$Category class$>$ list, making the outputs standardised and interpretable.
4.2.2 User Prompt
In this experiment, the User Prompt ($T_{U}$) combines both visual input and accompanying metadata, including {image_id} and {latitude, longitude}. These metadata elements help the vision–language models (VLMs) format their outputs more consistently and provide geographic context to guide interpretation.
Speed-related attributes such as “Speed Limit,” “Motorcycle Speed Limit,” and “Speed Management” are essential for iRAP assessments but are often not directly observable from imagery alone. To address this, our framework incorporates geographic metadata (e.g., GPS coordinates) into the VLM prompt, enabling the model to associate locations with region-specific traffic regulations learned from training data. For instance, rural roads in certain provinces may visually resemble one another but differ in legal speed limits— context, the model learns to infer through spatial priors. By combining visual input with location-aware prompts, VLMs draw on both scene appearance and broader regulatory patterns seen during pretraining. This multimodal design improves the model’s ability to classify speed-related attributes with greater accuracy and contextual relevance, especially when explicit signage is absent.
4.3 Vision Language Models
This work evaluates Gemini-1.5-flash and GPT-4o-mini as $f$ for their potential to replicate the work of road safety assessors under the iRAP standard. These models require no additional training or significant computational resources, making them accessible for use by local stakeholders. As demonstrated by Yue et al. [47], Gemini and GPT outperform other models in the Multi-discipline Multimodal Understanding and Reasoning benchmark. The experiments were conducted through the Gemini and the OpenAI APIs.
To assess their performance, we analysed 337 road segments (1348 images) with Gemini-1.5-flash and GPT-4o-mini, alongside ResNet and VGGNet baseline models. An additional set of 7 segments with unseen attributes was used to evaluate zero-shot classification capabilities.
4.4 Image Processor for Mapillary Imagery
Crowdsourced Street View Images (SVIs) are accessible on various platforms, with Mapillary being one of the most well-known, providing an API for image downloads. Hence, we evaluate the system using the V-RoAst framework with Gemini-1.5-flash to classify all 52 attributes. Then, we combine these results with additional information—including operating speed and ADDT from the ground truth—to determine the star rating outcomes.
For this work, images were obtained using a 50-metre buffer around ThaiRAP locations under the condition that the images were captured within one year of the collecting date. However, only 42 road segments were found to have corresponding Mapillary images, yielding 165 images. Panoramic images were converted to 1200x1600 binocular view images to align with the ThaiRAP data format.
It is important to note that some attributes may differ from the iRAP ground truth. For example, the number of vehicles shown in the image (observed flow) and the number of vehicles parked in the captured scene can vary. To validate the automatic use of V-RoAst, we did not verify whether the downloaded images originated from the exact same location, which could potentially affect the predicted star rating.
5 Experiments
5.1 Implementation Detail
Baseline
Our work used VGGNet and ResNet as baseline models to compare with our proposed approach. These models were selected because they have been widely used in previous work on road safety assessment and multi-task classification [40, 23, 24]. Since they are designed for single-task classification, we adapt the architectures for a multi-attribute coding problem, where a single encoder is shared across all tasks, and separate decoders are allocated for each individual task.
  • VGGNet [39] is a deep neural network model known for its simple architecture that stacks multiple convolutional layers with small 3x3 filters, achieving high performance in image classification tasks and becoming a widely used baseline in computer vision.
  • ResNet [13] is also a deep convolutional neural network that utilises residual blocks and skip connections to enhance feature learning at various abstraction levels, making it highly effective for image classification and transfer learning tasks.
Evaluation Metrics
We report results at two granularities. Overall performance is the macro-average of each metric across the iRAP attribute groups (“sections”), summarising model behaviour over broad feature types. Attribute-level performance is computed for each of the 52 individual attributes, revealing fine-grained strengths and weaknesses. When a road segment carries multiple risk classes, we follow iRAP convention and keep the first (highest-risk) label. For every attribute we calculate accuracy, precision, recall, and F1-score, then macro-average them so that attributes with few samples contribute equally to the final score.
5.2 Results and Discussions
5.2.1 Overall Performance
Table 1: Macro-averaged Accuracy, Precision, Recall, and F1-score for overall and unseen class performance. Bold highlights the best-performing model.
Group Attribute Model Acc Pre Rec F1
All Attributes CNN Models
   VGG 0.96 0.76 0.74 0.75
   ResNet 0.96 0.88 0.86 0.86
VLM Models
   GPT 0.75 0.46 0.47 0.42
   Gemini 0.82 0.49 0.50 0.47
Unseen Classes CNN Models
   VGG 0.24 0.18 0.23 0.18
   ResNet 0.29 0.39 0.25 0.21
VLM Models
   GPT 0.48 0.34 0.36 0.34
   Gemini 0.62 0.45 0.46 0.43
Table 1 presents the macro-averaged accuracy, precision, recall, and F1-score across all 52 attributes. Among the supervised baselines, ResNet achieves the highest macro performance with an accuracy of 0.96, precision of 0.88, recall of 0.86, and F1-score of 0.86. VGGNet also performs strongly, with slightly lower recall and precision, indicating solid but less consistent classification across attributes.
In contrast, the zero-shot VLMs show reduced performance. Gemini-1.5-flash achieves a macro-accuracy of 0.82 and an F1-score of 0.47, while GPT-4o-mini performs slightly behind. Despite this performance gap, VLMs offer a crucial advantage—the ability to predict all possible iRAP attribute classes without requiring task-specific training data. This is particularly valuable in data-sparse or label-scarce settings.
On the other hand, the Unseen Classes, those with highly imbalanced or missing training data (and thus excluded from the supervised baselines), VLMs excel. Gemini-1.5-flash achieves a macro-accuracy of 0.62 and an F1-score of 0.43, demonstrating its potential for generalised attribute classification without training.
Table 2: Macro-averaged Accuracy, Precision, Recall, and F1-score across iRAP-defined group attributes. Bold highlights the best-performing model.
Group Attribute Model Acc Pre Rec F1
Observed Flows CNN Models
   VGG 0.99 0.83 0.83 0.83
   ResNet 0.99 0.89 0.99 0.91
VLM Models
   GPT 0.95 0.47 0.49 0.47
   Gemini 0.98 0.68 0.65 0.66
Speed Limits CNN Models
   VGG 0.99 0.72 0.73 0.73
   ResNet 0.98 0.98 0.98 0.98
VLM Models
   GPT 0.52 0.46 0.52 0.34
   Gemini 0.84 0.55 0.53 0.53
Mid-block CNN Models
   VGG 0.97 0.67 0.64 0.64
   ResNet 0.98 0.92 0.88 0.90
VLM Models
   GPT 0.77 0.59 0.58 0.52
   Gemini 0.85 0.60 0.64 0.59
Roadside CNN Models
   VGG 0.90 0.72 0.73 0.72
   ResNet 0.91 0.76 0.72 0.74
VLM Models
   GPT 0.33 0.27 0.22 0.18
   Gemini 0.36 0.20 0.21 0.17
Intersections CNN Models
   VGG 0.95 0.76 0.73 0.74
   ResNet 0.97 0.82 0.77 0.79
VLM Models
   GPT 0.88 0.46 0.48 0.45
   Gemini 0.90 0.49 0.48 0.46
5.2.2 Performance by Attribute Group
Subgroup analyses highlight nuanced differences in Table 2. For spatially grounded attributes, such as roadside attributes, supervised models outperform VLMs significantly. Their training allows precise estimation of spatial and geometric features.
However, for attributes that are less spatially dependent, such as observed flows or mid-block attributes, VLMs show more competitive performance. Notably, Gemini-1.5-flash consistently outperforms GPT-4o-mini across all attribute groups, suggesting stronger generalisation and reasoning capabilities in visual question answering tasks.
5.2.3 Attribute-Level Analysis
Table 3 in the Appendix illustrates per-attribute performance. Baseline models were trained only on attributes with sufficient class diversity, excluding 11 attributes with a single class. VLMs are not constrained by this limitation and generate predictions for all attributes, offering broader applicability.
We observe that VLMs perform better when recognising prominent visual cues (e.g., presence of a school zone, number of lanes) but underperform in estimating distances or interpreting scene geometry. These results suggest that while VLMs are effective for attribute presence/absence tasks, they may struggle with precise spatial reasoning [36].
5.2.4 Qualitative Assessment using VQA
A key advantage of VLMs lies in their ability to perform VQA, enabling users–regardless of technical background–to iteratively refine model outputs through prompt adaptation. This makes VLMs particularly suitable for participatory or practitioner-led applications, where model interpretability and adaptability are crucial.
Figure 6 illustrates how prompt tuning can be employed to guide model reasoning and address misclassification. For example, in the bottom-right case, none of the four models tested correctly identified the “Sidewalk (passenger-side)” attribute. The ground truth label referred to an informal path located more than 1 metre from the main carriageway, a feature not clearly visible in the street-view image. This discrepancy highlights a common challenge that visual cues may be insufficient or ambiguous, requiring local contextual knowledge for accurate classification.
In such cases, supplementing street-level imagery with additional data sources, such as satellite views or local GIS layers, could help disambiguate difficult scenes. The interactive nature of VQA-based models also allows for on-the-fly prompt modifications to explore alternative interpretations, further enhancing usability for road safety assessments in resource-constrained environments.
Figure 6: Qualitative assessment of VLM performance using VQA with correct (green) and wrong (red) answers
5.2.5 Automatic Road Assessment
Figure 7 presents the confusion matrix for star rating predictions (motorcyclists) using Mapillary images with V-RoAst (Gemini-1.5-flash). The results demonstrate the model’s effectiveness in identifying high-risk roads with star ratings below 3, highlighted in the red box.
The flexibility of the V-RoAst framework lies in its integration of {local_context}, allowing local stakeholders to tailor prompts based on their expertise and validate results against ground truth data. This adaptability enables precise identification of high-risk areas, supporting road safety initiatives and investment prioritisation. Although Mapillary’s current coverage is limited, its crowdsourced platform serves as a valuable resource that transport authorities can leverage [41, 16]. Similar applications using commercial platforms like Google Street View may be explored, subject to licensing agreements.
Additional inputs, such as Annual Average Daily Traffic (AADT) and operating speeds, are essential to achieving comprehensive safety assessments. While this information is not derived directly from image labels, they are critical in determining road safety performance and star ratings.
Figure 7: Star rating (motorcyclists) confusion matrix of using crowdsourced imagery with V-RoAst and ground truth from ThaiRAP
6 Conclusion and Future work
This work introduced V-RoAst, a zero-shot prompting approach VLMs such as Gemini-1.5-flash and GPT-4o-mini for road safety assessment. Since VLMs enable flexible prompt engineering and results in potential adaptability improvements which is crucial for resource-limited settings in LMICs. VLMs excel at simple, interpretable attributes but struggle with spatially complex or implicit features like informal sidewalks. Limited street view coverage also impacts accuracy, highlighting the potential benefit of integrating satellite imagery for added context.
Despite current gaps, VLMs have the potential to complement traditional methods under the iRAP framework. Future research should focus on:
  • In-context learning: Applying few-shot or Chain-of-Thought prompting for better classification.
  • Fine-tuning: Adapting VLMs (e.g., with LoRA) to improve spatial and contextual understanding.
  • Multimodal integration: Combining street view with satellite and GIS data to fill visual gaps.
  • Cross-region validation: Testing VLMs in diverse environments for broader applicability.
  • Stakeholder usability: Creating accessible tools for local authorities with minimal technical expertise.
In summary, VLMs are not yet a full substitute for trained coders but could become a scalable solution to support global road safety, particularly in under-resourced regions, through continued refinement and integration.
Appendix
Table 3: Performance comparison across all iRAP-defined attributes using four models: VGG, Res(Net), GPT(-4o-mini), and Gem(ini-1.5-flash). Attribute group abbreviations: OF = Observed Flow, F = Facilities, DS = Driver-Side, PS = Passenger-Side, PC = Pedestrian Crossing, POF = Pedestrian Observed Flow, RS = Roadside Severity, SZ = School Zone. Metrics include Accuracy, Precision, Recall, and F1-score. For each attribute group, the last row (“All”) reports the macro-average across all attributes. Bold indicates the best-performing model.
Group Attribute Attribute Accuracy Precision Recall F1
CNN VLM CNN VLM CNN VLM CNN VLM
VGG Res GPT Gem VGG Res GPT Gem VGG Res GPT Gem VGG Res GPT Gem
Observed Flow Motorcycle (OF) 0.96 0.96 0.80 0.93 0.48 0.67 0.20 0.58 0.48 0.97 0.27 0.42 0.48 0.73 0.20 0.46
Bicycle (OF) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
(POF) across the road 1.00 1.00 0.50 1.00 0.50 1.00 0.50 1.00
(POF) along the road (DS) 1.00 1.00 0.99 0.99 1.00 1.00 0.33 0.33 1.00 1.00 0.33 0.33 1.00 1.00 0.33 0.33
(POF) along the road (PS) 1.00 1.00 0.99 0.99 1.00 1.00 0.33 0.50 1.00 1.00 0.33 0.50 1.00 1.00 0.33 0.50
All Observed Flow 0.99 0.99 0.95 0.98 0.83 0.89 0.47 0.68 0.83 0.99 0.49 0.65 0.83 0.91 0.47 0.66
Speed Limits Speed limit 0.99 0.98 0.44 0.74 0.65 0.98 0.34 0.50 0.67 0.97 0.33 0.48 0.66 0.98 0.30 0.48
Motorcycle speed limit 1.00 1.00 0.54 0.96 0.48 1.00 0.55 0.86 0.50 1.00 0.76 0.73 0.49 1.00 0.44 0.78
Truck speed limit 1.00 1.00 0.62 0.94 0.48 1.00 0.38 0.57 0.50 1.00 0.52 0.48 0.49 1.00 0.34 0.52
Differential speed limits 0.99 0.98 0.21 0.74 1.00 0.97 0.60 0.39 1.00 0.96 0.51 0.46 1.00 0.97 0.19 0.43
Speed management 0.98 0.96 0.80 0.80 1.00 0.97 0.40 0.40 1.00 0.96 0.50 0.50 1.00 0.97 0.45 0.45
All Speed Limits 0.99 0.98 0.52 0.84 0.72 0.98 0.46 0.55 0.73 0.98 0.52 0.53 0.73 0.98 0.34 0.53
Mid-block Number of lanes 0.95 0.95 0.14 0.55 0.66 0.82 0.29 0.30 0.44 0.82 0.32 0.54 0.49 0.82 0.09 0.26
Lane width 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Curvature 0.99 0.99 0.98 0.98 0.50 1.00 0.49 0.62 0.48 1.00 0.50 0.62 0.49 1.00 0.49 0.62
Quality of curve 0.99 0.99 0.98 0.98 0.50 1.00 0.49 0.62 0.48 1.00 0.50 0.62 0.49 1.00 0.49 0.62
Upgrade cost 0.93 0.96 0.49 0.61 0.92 0.98 0.57 0.39 0.83 0.97 0.34 0.45 0.86 0.97 0.25 0.40
Median type 0.93 0.95 0.19 0.72 0.67 0.97 0.17 0.21 0.70 0.84 0.22 0.25 0.68 0.89 0.08 0.23
Skid resistance / grip 0.98 0.99 0.80 0.80 0.98 0.98 0.40 0.73 0.95 0.95 0.50 0.51 0.96 0.96 0.44 0.47
Road condition 0.00 0.00 0.99 0.98 0.50 0.50 0.50 0.49 0.50 0.49
Vehicle parking 0.99 1.00 0.79 0.97 0.48 0.48 0.33 0.50 0.50 0.50 0.27 0.49 0.49 0.49 0.30 0.49
Grade 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Roadworks 1.00 0.99 0.50 0.50 0.50 0.50 0.50 0.50
Sight distance 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Delineation 1.00 1.00 1.00 1.00 0.50 1.00 0.50 0.50 0.48 1.00 0.50 0.50 0.49 1.00 0.50 0.50
Street lighting 0.96 0.97 0.89 0.94 1.00 1.00 0.62 0.75 1.00 1.00 0.68 0.72 1.00 1.00 0.64 0.73
Service road 0.97 0.98 0.12 0.10 0.47 0.98 0.55 0.05 0.50 0.75 0.51 0.50 0.48 0.83 0.11 0.09
Centreline rumble strips 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All Mid-block 0.97 0.98 0.77 0.85 0.67 0.92 0.59 0.60 0.64 0.88 0.58 0.64 0.64 0.90 0.52 0.59
Roadside (RS) - (DS) distance 0.97 0.96 0.13 0.00 0.91 0.94 0.25 0.00 0.91 0.98 0.17 0.00 0.91 0.96 0.10 0.00
(RS) - (DS) object 0.93 0.94 0.25 0.82 0.46 0.46 0.18 0.27 0.48 0.48 0.10 0.29 0.47 0.47 0.10 0.28
(RS) - (PS) distance 0.86 0.88 0.36 0.04 0.60 0.66 0.39 0.03 0.58 0.52 0.38 0.26 0.57 0.55 0.22 0.04
(RS) - (PS) object 0.71 0.76 0.24 0.26 0.82 0.89 0.29 0.06 0.80 0.87 0.26 0.16 0.80 0.88 0.16 0.09
Shoulder rumble strips 0.99 0.98 0.99 0.96 0.48 0.48 0.50 0.50 0.50 0.50 0.50 0.48 0.49 0.49 0.50 0.49
Paved shoulder - (DS) 0.98 0.96 0.11 0.02 0.91 0.88 0.22 0.25 0.91 0.84 0.03 0.01 0.91 0.86 0.06 0.01
Paved shoulder - (PS) 0.87 0.89 0.19 0.41 0.86 0.98 0.09 0.31 0.87 0.94 0.11 0.27 0.86 0.96 0.10 0.25
All Roadside 0.90 0.91 0.33 0.36 0.72 0.76 0.27 0.20 0.72 0.73 0.22 0.21 0.72 0.74 0.18 0.17
Intersections Land use - (DS) 1.00 1.00 0.38 0.99 1.00 1.00 0.33 0.33 1.00 1.00 0.46 0.33 1.00 1.00 0.19 0.33
Land use - (PS) 0.86 0.88 0.70 0.26 0.55 0.55 0.26 0.21 0.53 0.54 0.31 0.17 0.54 0.54 0.27 0.11
Area type 0.96 0.98 0.90 0.82 1.00 1.00 0.68 0.66 1.00 1.00 0.58 0.86 1.00 1.00 0.61 0.68
(PC) (F) - inspected road 0.94 0.96 0.94 0.95 0.48 0.48 0.31 0.57 0.48 0.47 0.33 0.39 0.48 0.48 0.32 0.42
(PC) (F) - intersecting road 0.99 0.99 0.99 0.99 1.00 1.00 0.33 0.33 1.00 1.00 0.33 0.33 1.00 1.00 0.33 0.33
(PC) quality 0.95 0.95 0.93 0.93 0.48 0.48 0.47 0.47 0.48 0.47 0.50 0.50 0.48 0.48 0.48 0.48
Pedestrian fencing 1.00 1.00 1.00 1.00 1.00 1.00 0.50 0.50 1.00 1.00 0.50 0.50 1.00 1.00 0.50 0.50
Sidewalk - (DS) 0.89 0.98 0.50 0.33 0.45 0.33 0.47 0.33
Sidewalk - (PS) 0.92 0.93 0.77 0.83 0.34 0.72 0.18 0.31 0.32 0.74 0.21 0.20 0.33 0.73 0.20 0.20
(F) for bicycles 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
(F) for motorised two wheelers 1.00 1.00 0.50 1.00 0.50 1.00 0.50 1.00
(SZ) crossing supervisor 0.96 0.98 0.96 0.96 0.73 0.98 0.48 0.32 0.73 0.75 0.50 0.33 0.73 0.83 0.49 0.33
(SZ) warning 0.96 0.98 0.96 0.96 0.98 0.98 0.48 0.32 0.75 0.75 0.50 0.33 0.83 0.83 0.49 0.33
All Intersections 0.95 0.97 0.88 0.90 0.76 0.82 0.46 0.49 0.73 0.77 0.48 0.48 0.74 0.79 0.45 0.46